Grade 12Physical Chemistry

The volume of co2 obtained at stp by the complete thermal decomposition of 25 g of caco3 is

Pawan Prajapati
5 Years agoGrade 12

1 Answer

abhay gupta
5 Years ago
Answer:

CaCO3 molar mass is 100 u.

or it can be represented as 100 grams when it is expressed as molecular mass.

CaCO3 ---------> CaO + CO2

CO2 molar mass is 44 u, it's molecular mass is 44 grams.

This implies 100 g of CaCO3 gives 44 g of CO2.

Therefore 25 g of CaCO3 gives 44/100 * 25 = 11 grams.

One mole of CO2 is 44 grams.

Therefore in 11 grams of CO2 , there would be 1/44 * 11 = 0.25 moles.

one mole of CO2 occupies 22.4 L of volume, therefore 0.25 moles of CO2 has a volume of 22.4 * 0.25 =

5.6 L of volume.

thank you !